drug a (t1/2= 18 hours, v= 80l fe= 0.2 fu= 0.1) is administered as multiple 200 mg iv bolus every 12 hours. predict max concentration at steady state.
Answer #1

Drug A

Half Life - 18hours

V - 80L

fe - 0.2

fu - 0.1

Drug A is administered as multiple 200mg IV bolus every 12 hours .

At steady state, the amount of drug lost in each interval equals the amount gained, that is the dose multiplied by the bioavailability.

steady state concentration can be assessed using the equation:

C​​​​​​ssav = F×dose/ CL ×dosing interval

τ = dosing interval, hr = 12 hourly

Dose = 200mg

F = Bioavailability

CL = Clearance

CL = 0.693/ half life × Vd

= 0.693/ 18 × 80

= 3.08L /hr (total clearance ) = 3080ml /hr

F = 100% because drug is given IV

Steady state concentration = 100× 200mg / 3080ml /hr × 12 hourly

= 0.54mg / ml